OpenStudy (anonymous):

2B twelve + 3A twelve

sammixboo (sammixboo):

Do you have anymore info?

OpenStudy (lizz123):

\[\frac{ 2b }{ 12 }+\frac{ 3a }{ 12 }=\]

OpenStudy (lizz123):

@verodisque

OpenStudy (anonymous):

Is arithmetic in different base

OpenStudy (anonymous):

Any one knows how to solve different base 2B twelve + 3A twelve

OpenStudy (anonymous):

\[3 A+2 B=35_{10}+46_{10}=81_{10}=69_{12} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

B1 twelve - 57 twelve?